OpenStudy (anonymous):

help prove that (sinx cotx+cosx)/2cotx=sinx ??

OpenStudy (amoodarya):

write cot = cos /sin then simplify

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

sinx cosx/sinx+cosx=2cosx 2cosx/2cosx/sinx=sinx

OpenStudy (anonymous):

See. There are many ways to solve it. 1. take the primary formulae for every trigonometric identities. i.e. sinx=p/h, cotx=b/p cosx=b/h and all that then you may sove it. but its a li'l bit leanthy. 2. given (sinx.cotx+cosx)/2cotx =LHS put cotx=cosx/sonx so , (sinx(cosx/sinx)+cosx)/2cotx = (cosx+cosx)/2cotx =2cosx/2cotx =cosx/(cosx/sinx) =sinx =RHS ; hance proved :)

OpenStudy (amoodarya):

\[\frac{ \sin x \cot x +\cos x }{ 2\cot x}=\frac{ \sin x \frac{ \cos x }{ } +\cos x }{ 2\frac{ \cos x }{ \sin x }}=\\ \frac{ 2\cos x }{ \ }/\frac{ 2\cos x }{ \sin x }\]